OpenStudy (anonymous):

MEDALS Olivia needs $36 to buy a coat. She has saved $8 and plans to work as a babysitter to earn $7 per hour. Which inequality shows the minimum number of hours, n, that Olivia should work as a babysitter to earn enough to buy the coat?

OpenStudy (anonymous):

8 + 7n ≤ 36, so n ≤ 4 8 + 7n ≥ 36, so n ≥ 4 7n ≥ 36 + 8, so n ≥ 6.3 7n ≤ 36 + 8, so n ≤ 6.3

OpenStudy (anonymous):

c is wrong

OpenStudy (anonymous):

Olivia needs to find the discount rack.

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

or walk through it with me

OpenStudy (anonymous):

A, Because you already have 8 so that would be added, so it would be 8 +. This narrows the answer down to A and B. if you earn 7 dollars per hour, \[7*4=28 +8=36\]

OpenStudy (anonymous):

thanks